OpenStudy (anonymous):

What are the solutions of the system? y=x^2+4x-5 y=2x-2 A. (–3, –8) and (1, 0) B. (–3, –8) and (–1, –4) C. (–3, 0) and (1, –8) D. no solution

OpenStudy (anonymous):

so one please help?

OpenStudy (anonymous):

I think it is D but not sure.

OpenStudy (callisto):

y=x^2+4x-5 -(1) y=2x-2 -(2) (1)- (2): 0 = x^2 +4x -5 - (2x-2) 0 = x^2 +2x -3 (x+3)(x-1) =0 x= -3 or x =1 When x = -3. y = (-3)^2 + 4(-3) -5 = -8 So, (-3, -8) is a pair of solutions When x = 1 y = (1)^2 + 4(1) -5 = 0 So, (1, 0) is another pair of solutions.

OpenStudy (anonymous):

so the answer would be A.

OpenStudy (callisto):

Yes... I hope you know where you got wrong...

OpenStudy (anonymous):

i do thanks

OpenStudy (callisto):

Welcome :)